Classify the Critical points of f(x,y)=4+x^3+y^3-3xy


1
Expert's answer
2022-03-01T13:55:04-0500

"f'_x=3x^2-3y"

"f'_y=3y^2-3x"

Find the critical point(s)


"3x^2-3y=0""3y^2-3x=0"

"y=x^2""x^4-x=0"

"x(x-1)(x^2+x+1)=0""x=y^2"

Critical point "(0, 0)," critical point "(1, 1)."


Use Second Derivatives Test


"f''_{xx}=6x"

"f''_{xy}=-3=f''_{yx}"

"f''_{yy}=6y"

"D=\begin{vmatrix}
 f''_{xx}& f''_{xy} \\
 f''_{yx} & f''_{yy}
\end{vmatrix}=\begin{vmatrix}
 6x & -3 \\
 -3 & 6y
\end{vmatrix}=36xy-9"

Critical point "(0, 0)"

"D(0, 0)=0-9=-9<0"

"f(0,0)" is not a local maximum or minimum.

Point "(0,0)" is a saddle point of "f."


Critical point "(1, 1)"

"D(1, 1)=36(1)(1)-9=27>0"

"f''_{xx}(1,1)=6(1)=6>0."


"f(1,1)" is a local minimum.